# Protein Structure Tokenization via Geometric Byte Pair Encoding

### v1

@@ Appendix A Ablation Studies
Figure 5: We ran an ablation for GeoBPE version with $|\mathcal{V}|=600$, keeping all parameters the same but toggling whether glue opt is skipped in Alg. 16. We visualize the original (center), GeoBPE (left) and GeoBPE without glue opt (right) backbone states.
Rigid body refinment as an essential step for preserving fold integrity. If we omit the glue optimization from Algs 16 and 9 altogether, we see the effects in Fig. 5. For that experiment, we find avg. RMSD increase $1.66\rightarrow 4.39$, and avg. LDDT drop $0.73\rightarrow 0.69$ when glue opt is turned off. Rigid body refinement preserves the overall fold and modular architecture; turning it off causes individual domains to distort – the parallel strands drift apart – as well as the overall configuration to lose its integrity. Over the course of many time steps, global drift accumulate as local rounding occurs. Rigid body refinement is an indispensable subroutine for ensuring the overall quantization faithfully reproduces the fold integrity.
Increasing $M_{\text{max}}$ beyond a certain threshold does not yield additional distortion benefits. We did a study comparing GeoBPE ($\mathcal{V}=6000$, $M_{\text{max}}\leftarrow 5000$, full settings in App. I) with “higher-resolution” settings $\texttt{bins}\leftarrow\{1:5000\}$, $M_{\text{max}}\leftarrow 20000$. Interestingly, we found overall RMSD/LDDT did not improve ($1.40$ vs $1.39$, $0.76$ vs $0.75$, both in favor of the incumbent) despite increased computational expenditure spent on Alg. 8. The most likely explanation is there is no marginal utility increasing $M_{\text{max}}$ beyond $5000$, and differences in distortion rates are likely due to the numerical stability of Alg. 10 more so than the hyperparameters.
@@ Appendix C Additional Results
Table 4 contains additional tasks Rep, Ept, FlexRMSF, and FlexBFactor. Table 5 contains additional task data Repeat, Ept, Atlas, Homo. Task abbrevations were defined in App. B.
@@
ProteinMPNN MIF FoldSeek ProTokens ESM3 VanillaVQ AminoAseed GeoBPE (v.s. ESM3)
@@
[…]
@@ Appendix D Token Efficiency Metrics
Table 6: We evaluate token efficiency of GeoBPE across varying $|\mathcal{V}|\in\{600,2500,6000\}$, as reported in Figure 3.
@@
Table 7: We adopt the Small Structure Language Model evaluation protocol described in App. E. We sample 100 PDB structures.
@@
[…]
@@ Appendix E Small Structure Language Model Evaluation
This section specifies the protocol used to evaluate generative efficiency via training a small decoder-only LM over the joint geometric vocabulary learned by GeoBPE. It incorporates the mask constraints used at generation time (Alg. 4) during training to ensure consistency between training and sampling. The same procedure is used for evaluating VQ-VAEs. For training and sampling, the only difference is dropping the mask constraints. For inference, the samples are passed through the VQ-VAE decoder to construct backbone coordinates instead of assembling the backbone directly via Alg. 5. This required considerably more resources, and we discuss how we implemented this in App. E.6.
@@
0: Corpus of proteins $\{t_{\tau}\}_{\tau=1}^{T}$ with final segmentations $\{\mathcal{P}^{(\tau)}\}$ and assigned medoids; joint vocabulary $\Sigma$ and tokenizers from Alg. 14, 15; a decoder-only Transformer $\mathsf{Tr}_{\theta}:\Sigma^{\ast}\!\to\!\Delta^{|\Sigma|}$ with causal mask; special BOS/EOS (optional); training steps $S$, optimizer $\mathcal{O}$.
@@
1:Dataset construction. For each $\tau$, build $x^{(\tau)}=\textsc{BackboneToSequence}(t_{\tau})$ (Alg. 15). Let $L_{\tau}=|x^{(\tau)}|$.
@@
0: Training corpus of tokenized backbones $\{x^{(\tau)}=(x^{(\tau)}_{1},\ldots,x^{(\tau)}_{L_{\tau}})\}_{\tau=1}^{T}$ constructed by Alg. 15 (motif, then $\theta,\omega,\phi$, repeating); valid sequence lengths satisfy $L_{\tau}\equiv 1\ (\mathrm{mod}\ 4)$ and end in a terminating motif token.
@@
0: Trained decoder-only Transformer $\mathsf{Tr}_{\theta}$ with vocabulary $\Sigma$ from Alg. 14; id blocks

### v2

@@ Appendix A Ablation Studies
Figure 6: We ran an ablation for GeoBPE version with $|\mathcal{V}|=600$, keeping all parameters the same but toggling whether glue opt is skipped in Alg. 18. We visualize the original (center), GeoBPE (left) and GeoBPE without glue opt (right) backbone states.
Rigid body refinment as an essential step for preserving fold integrity. If we omit the glue optimization from Algs 18 and 9 altogether, we see the effects in Fig. 6. For that experiment, we find avg. RMSD increase $1.66\rightarrow 4.39$, and avg. LDDT drop $0.73\rightarrow 0.69$ when glue opt is turned off. Rigid body refinement preserves the overall fold and modular architecture; turning it off causes individual domains to distort – the parallel strands drift apart – as well as the overall configuration to lose its integrity. Over the course of many time steps, global drift accumulate as local rounding occurs. Rigid body refinement is an indispensable subroutine for ensuring the overall quantization faithfully reproduces the fold integrity.
Increasing $M_{\text{max}}$ beyond a certain threshold does not yield additional distortion benefits. We did a study comparing GeoBPE ($\mathcal{V}=6000$, $M_{\text{max}}\leftarrow 5000$, full settings in App. K) with “higher-resolution” settings $\texttt{bins}\leftarrow\{1:5000\}$, $M_{\text{max}}\leftarrow 20000$. Interestingly, we found overall RMSD/LDDT did not improve ($1.40$ vs $1.39$, $0.76$ vs $0.75$, both in favor of the incumbent) despite increased computational expenditure spent on Alg. 8. The most likely explanation is there is no marginal utility increasing $M_{\text{max}}$ beyond $5000$, and differences in distortion rates are likely due to the numerical stability of Alg. 12 more so than the hyperparameters.
@@ Appendix C Additional Results
ProteinMPNN MIF FoldSeek ProTokens ESM3 VanillaVQ AminoAseed GeoBPE-Transfer (v.s. ESM3)
@@
FlexBFactor Fold 31.88 34.60 4.17 6.67 23.60 22.32 21.30 37.28 (+57.97%)
SupFam 34.56 35.23 6.99 5.47 25.80 23.73 21.76 35.61 (+38.02%)
FlexNEQ Fold 69.69 65.32 5.71 12.98 45.05 35.95 49.64 56.65 (+25.75%)
SupFam 68.69 64.82 2.66 10.51 35.45 35.61 50.15 53.98 (+52.27%)
@@
Main Text Tables. Table 6 contains additional tasks Rep, Ept, FlexRMSF, and FlexBFactor. Table 7 contains additional task data Repeat, Ept, Atlas, Homo. Task abbrevations are defined in App. B.
Table 8: Secondary structure element (SSE) agreement results. Setup is the same as in Sec. G, but stratified over 8 basic secondary structure building blocks. Annotations are obtained from DSSP.
Metric / SSE BindInt BindBio BindShake CatInt CatBio Con Repeat Ept Atlas Homo
Mean recall H 29.11 (98.22) 33.13 (97.42) 30.76 (96.39) 33.97 (98.1) 32.93 (96.48) 29.11 (98.22) 33.13 (97.42) 30.76 (96.39) 33.97 (98.1) 32.93 (96.48)
Mean recall G 19.15 (99.4) 31.13 (98.61) 30.87 (98.55) 34.66 (98.75) 32.29 (98.71) 19.15 (99.4) 31.13 (98.61) 30.87 (98.55) 34.66 (98.75) 32.29 (98.71)
Mean recall I 10.42 (100.0) 30.75 (100.0) 17.59 (100.0) 10.85 (99.26) 30.85 (99.68) 10.42 (100.0) 30.75 (100.0) 17.59 (100.0) 10.85 (99.26) 30.85 (99.68)
Mean recall E 41.24 (99.56) 34.75 (98.46) 36.17 (97.49) 30.82 (97.51) 34.93 (97.99) 41.24 (99.56) 34.75 (98.46) 36.17 (97.49) 30.82 (97.51) 34.93 (97.99)
Mean recall B 38.13 (100.0) 32.24 (100.0) 38.72 (100.0) 40.22 (100.0) 31.67 (100.0) 38.13 (100.0) 32.24 (100.0) 38.72 (100.0) 40.22 (100.0) 31.67 (100.0)
Mean recall T 32.78 (98.74) 33.08 (97.4) 32.26 (96.41) 33.75 (96.67) 32.91 (97.08) 32.78 (98.74) 33.08 (97.4) 32.26 (96.41) 33.75 (96.67) 32.91 (97.08)
Mean recall S 33.92 (99.32) 33.29 (98.0) 32.91 (97.72) 36.01 (97.4) 33.12 (97.96) 33.92 (99.32) 33.29 (98.0) 32.91 (97.72) 36.01 (97.4) 33.12 (97.96)
Mean recall - 32.7 (97.81) 32.76 (97.17) 33.02 (96.16) 34.71 (96.86) 32.65 (96.03) 32.7 (97.81) 32.76 (97.17) 33.02 (96.16) 34.71 (96.86) 32.65 (96.03)
Mean precision H 31.34 (54.95) 34.06 (48.88) 31.33 (51.22) 35.54 (48.86) 33.32 (46.04) 31.34 (54.95) 34.06 (48.88) 31.33 (51.22) 35.54 (48.86) 33.32 (46.04)
Mean precision G 17.93 (79.42) 29.28 (65.93) 28.95 (62.87) 32.3 (65.07) 30.32 (64.77) 17.93 (79.42) 29.28 (65.93) 28.95 (62.87) 32.3 (65.07) 30.32 (64.77)
Mean precision I 10.09 (94.67) 29.93 (87.7) 16.55 (87.53) 10.37 (93.27) 30.16 (88.4) 10.09 (94.67) 29.93 (87.7) 16.55 (87.53) 10.37 (93.27) 30.16 (88.4)
Mean precision E 39.54 (47.86) 33.59 (46.21) 34.84 (45.6) 29.58 (53.03) 33.53 (45.56) 39.54 (47.86) 33.59 (46.21) 34.84 (45.6) 29.58 (53.03) 33.53 (45.56)
Mean precision B 26.8 (74.03) 21.86 (71.01) 26.66 (68.64) 26.21 (61.46) 21.2 (69.63) 26.8 (74.03) 21.86 (71.01) 26.66 (68.64) 26.21 (61.46) 21.2 (69.63)
Mean precision T 28.59 (53.66) 28.55 (50.1) 27.98 (50.94) 29.18 (52.52) 28.13 (48.36) 28.59 (53.66) 28.55 (50.1) 27.98 (50.94) 29.18 (52.52) 28.13 (48.36)
Mean precision S 26.78 (56.59) 25.28 (50.82) 24.66 (49.22) 26.52 (45.95) 25.0 (50.96) 26.78 (56.59) 25.28 (50.82) 24.66 (49.22) 26.52 (45.95) 25.0 (50.96)
Mean precision - 27.07 (54.17) 26.5 (48.75) 26.32 (48.44) 28.07 (51.65) 25.99 (48.15) 27.07 (54.17) 26.5 (48.75) 26.32 (48.44) 28.07 (51.65) 25.99 (48.15)
Mean f1 H 29.42 (61.9) 33.12 (60.16) 30.7 (62.58) 34.0 (59.45) 32.79 (58.18) 29.42 (61.9) 33.12 (60.16) 30.7 (62.58) 34.0 (59.45) 32.79 (58.18)
Mean f1 G 18.44 (83.1) 30.03 (72.52) 29.75 (70.88) 33.34 (72.73) 31.15 (72.15) 18.44 (83.1) 30.03 (72.52) 29.75 (70.88) 33.34 (72.73) 31.15 (72.15)
Mean f1 I 10.25 (95.59) 30.24 (89.98) 17.04 (89.31) 10.59 (94.02) 30.38 (90.53) 10.25 (95.59) 30.24 (89.98) 17.04 (89.31) 10.59 (94.02) 30.38 (90.53)
Mean f1 E 40.21 (59.65) 33.88 (60.28) 35.25 (59.17) 29.98 (63.68) 33.99 (60.01) 40.21 (59.65) 33.88 (60.28) 35.25 (59.17) 29.98 (63.68) 33.99 (60.01)
Mean f1 B 30.57 (75.28) 25.3 (73.05) 30.65 (71.04) 30.86 (64.57) 24.67 (71.8) 30.57 (75.28) 25.3 (73.05) 30.65 (71.04) 30.86 (64.57) 24.67 (71.8)
Mean f1 T 30.22 (62.79) 30.27 (61.7) 29.61 (63.41) 30.93 (64.04) 29.95 (61.29) 30.22 (62.79) 30.27 (61.7) 29.61 (63.41) 30.93 (64.04) 29.95 (61.29)
Mean f1 S 29.24 (61.77) 28.04 (58.79) 27.46 (57.52) 29.76 (54.36) 27.79 (59.56) 29.24 (61.77) 28.04 (58.79) 27.46 (57.52) 29.76 (54.36) 27.79 (59.56)
Mean f1 - 28.82 (61.51) 28.59 (60.08) 28.57 (59.89) 30.35 (63.16) 28.25 (60.75) 28.82 (61.51) 28.59 (60.08) 28.57 (59.89) 30.35 (63.16) 28.25 (60.75)
Mean iou H 28.4 (61.54) 32.22 (59.93) 29.94 (62.38) 32.95 (59.09) 31.96 (57.96) 28.4 (61.54) 32.22 (59.93) 29.94 (62.38) 32.95 (59.09) 31.96 (57.96)
Mean iou G 17.75 (83.07) 28.87 (72.45) 28.59 (70.78) 31.99 (72.63) 29.98 (72.11) 17.75 (83.07) 28.87 (72.45) 28.59 (70.78) 31.99 (72.63) 29.98 (72.11)
Mean iou I 10.09 (95.59) 29.58 (89.98) 16.52 (89.31) 10.31 (94.02) 29.67 (90.51) 10.09 (95.59) 29.58 (89.98) 16.52 (89.31) 10.31 (94.02) 29.67 (90.51)
Mean iou E 39.1 (59.46) 32.76 (59.84) 34.14 (58.81) 29.02 (63.37) 32.9 (59.53) 39.1 (59.46) 32.76 (59.84) 34.14 (58.81) 29.02 (63.37) 32.9 (59.53)
Mean iou B 26.79 (74.03) 21.86 (71.14) 26.66 (68.8) 26.21 (61.53) 21.2 (69.74) 26.79 (74.03) 21.86 (71.14) 26.66 (68.8) 26.21 (61.53) 21.2 (69.74)
Mean iou T 28.36 (61.64) 28.26 (59.8) 27.69 (61.71) 28.9 (62.33) 27.92 (59.16) 28.36 (61.64) 28.26 (59.8) 27.69 (61.71) 28.9 (62.33) 27.92 (59.16)
Mean iou S 26.63 (59.27) 25.15 (53.97) 24.47 (52.46) 26.41 (48.85) 24.9 (54.31) 26.63 (59.27) 25.15 (53.97) 24.47 (52.46) 26.41 (48.85) 24.9 (54.31)
Mean iou - 26.17 (57.07) 25.93 (54.13) 25.83 (53.73) 27.63 (57.52) 25.56 (53.91) 26.17 (57.07) 25.93 (54.13) 25.83 (53.73) 27.63 (57.52) 25.56 (53.91)
Segment recall H 29.35 (99.64) 33.14 (99.88) 30.94 (99.93) 33.77 (99.65) 33.07 (99.91) 29.35 (99.64) 33.14 (99.88) 30.94 (99.93) 33.77 (99.65) 33.07 (99.91)
Segment recall G 19.18 (100.0) 31.2 (100.0) 30.93 (100.0) 34.89 (100.0) 32.37 (99.85) 19.18 (100.0) 31.2 (100.0) 30.93 (100.0) 34.89 (100.0) 32.37 (99.85)
Segment recall I 10.42 (100.0) 30.79 (100.0) 17.63 (100.0) 10.91 (100.0) 30.94 (100.0) 10.42 (100.0) 30.79 (100.0) 17.63 (100.0) 10.91 (100.0) 30.94 (100.0)
Segment recall E 41.28 (100.0) 34.8 (99.36) 36.21 (99.8) 30.84 (100.0) 35.0 (99.68) 41.28 (100.0) 34.8 (99.36) 36.21 (99.8) 30.84 (100.0) 35.0 (99.68)
Segment recall B 38.01 (99.69) 31.7 (98.92) 38.05 (98.75) 39.75 (99.02) 31.2 (98.87) 38.01 (99.69) 31.7 (98.92) 38.05 (98.75) 39.75 (99.02) 31.2 (98.87)
Segment recall T 32.81 (99.4) 33.09 (99.07) 32.25 (98.3) 33.74 (98.46) 32.86 (98.61) 32.81 (99.4) 33.09 (99.07) 32.25 (98.3) 33.74 (98.46) 32.86 (98.61)
Segment recall S 33.73 (99.03) 33.02 (98.7) 32.5 (98.17) 35.39 (97.75) 32.77 (98.16) 33.73 (99.03) 33.02 (98.7) 32.5 (98.17) 35.39 (97.75) 32.77 (98.16)
Segment recall - 33.29 (98.57) 32.94 (98.35) 33.07 (98.3) 34.87 (98.83) 32.72 (98.52) 33.29 (98.57) 32.94 (98.35) 33.07 (98.3) 34.87 (98.83) 32.72 (98.52)
Segment precision H 39.24 (63.48) 34.23 (61.03) 34.32 (61.08) 29.43 (59.98) 33.02 (59.47) 39.24 (63.48) 34.23 (61.03) 34.32 (61.08) 29.43 (59.98) 33.02 (59.47)
Segment precision G 2.64 (82.35) 4.22 (72.63) 4.4 (71.91) 4.56 (75.06) 4.16 (70.14) 2.64 (82.35) 4.22 (72.63) 4.4 (71.91) 4.56 (75.06) 4.16 (70.14)
Segment precision I 1.96 (93.94) 2.25 (83.44) 1.2 (90.89) 0.68 (93.18) 1.68 (82.77) 1.96 (93.94) 2.25 (83.44) 1.2 (90.89) 0.68 (93.18) 1.68 (82.77)
Segment precision E 34.7 (62.4) 23.29 (64.43) 23.79 (64.45) 17.3 (65.59) 20.66 (63.88) 34.7 (62.4) 23.29 (64.43) 23.79 (64.45) 17.3 (65.59) 20.66 (63.88)
Segment precision B 1.28 (99.67) 1.08 (98.39) 1.18 (98.47) 1.2 (98.86) 0.96 (98.46) 1.28 (99.67) 1.08 (98.39) 1.18 (98.47) 1.2 (98.86) 0.96 (98.46)
Segment precision T 9.58 (70.95) 9.36 (72.49) 9.05 (71.02) 9.59 (71.11) 9.21 (72.8) 9.58 (70.95) 9.36 (72.49) 9.05 (71.02) 9.59 (71.11) 9.21 (72.8)
Segment precision S 8.12 (78.71) 6.83 (83.39) 6.37 (82.98) 7.36 (83.94) 6.53 (83.21) 8.12 (78.71) 6.83 (83.39) 6.37 (82.98) 7.36 (83.94) 6.53 (83.21)
Segment precision - 15.47 (81.35) 16.31 (81.4) 15.15 (80.22) 16.62 (79.37) 15.71 (83.16) 15.47 (81.35) 16.31 (81.4) 15.15 (80.22) 16.62 (79.37) 15.71 (83.16)
Segment f1 H 30.95 (63.48) 32.18 (61.08) 31.04 (61.09) 30.71 (60.02) 32.09 (59.54) 30.95 (63.48) 32.18 (61.08) 31.04 (61.09) 30.71 (60.02) 32.09 (59.54)
Segment f1 G 4.54 (82.35) 7.27 (72.63) 7.56 (71.91) 7.96 (75.06) 7.24 (70.1) 4.54 (82.35) 7.27 (72.63) 7.56 (71.91) 7.96 (75.06) 7.24 (70.1)
Segment f1 I 3.17 (93.94) 4.06 (83.44) 2.2 (90.89) 1.28 (93.18) 3.15 (82.77) 3.17 (93.94) 4.06 (83.44) 2.2 (90.89) 1.28 (93.18) 3.15 (82.77)
Segment f1 E 35.43 (62.51) 26.32 (64.46) 27.09 (64.47) 21.22 (65.66) 24.68 (63.9) 35.43 (62.51) 26.32 (64.46) 27.09 (64.47) 21.22 (65.66) 24.68 (63.9)
Segment f1 B 2.46 (99.67) 2.06 (98.38) 2.25 (98.44) 2.31 (98.86) 1.85 (98.46) 2.46 (99.67) 2.06 (98.38) 2.25 (98.44) 2.31 (98.86) 1.85 (98.46)
Segment f1 T 14.37 (71.22) 14.33 (73.02) 13.94 (71.35) 14.72 (71.68) 14.18 (73.37) 14.37 (71.22) 14.33 (73.02) 13.94 (71.35) 14.72 (71.68) 14.18 (73.37)
Segment f1 S 12.6 (79.62) 10.96 (83.99) 10.36 (83.6) 11.91 (84.44) 10.68 (83.93) 12.6 (79.62) 10.96 (83.99) 10.36 (83.6) 11.91 (84.44) 10.68 (83.93)
Segment f1 - 20.45 (84.94) 21.18 (86.73) 20.3 (86.08) 22.14 (85.95) 20.82 (88.68) 20.45 (84.94) 21.18 (86.73) 20.3 (86.08) 22.14 (85.95) 20.82 (88.68)
Table 9: Table 8 with averages over 10 datasets. We also report global averages over all 8 SSEs.
Metric H G I E B T S - Avg (HGIEBTS-)
Mean recall 31.98 (97.32) 29.62 (98.80) 20.09 (99.79) 35.58 (98.20) 36.20 (100.00) 32.96 (97.26) 33.85 (98.08) 33.17 (96.81) 31.68 (98.28)
Mean precision 33.12 (49.99) 27.76 (67.61) 19.42 (90.31) 34.22 (47.65) 24.55 (68.95) 28.49 (51.12) 25.65 (50.71) 26.79 (50.23) 27.50 (59.57)
Mean f1 32.01 (60.45) 28.54 (74.28) 19.70 (91.89) 34.66 (60.56) 28.41 (71.15) 30.20 (62.65) 28.46 (58.40) 28.92 (61.08) 28.86 (67.56)
Mean iou 31.09 (60.18) 27.44 (74.21) 19.23 (91.88) 33.58 (60.20) 24.54 (69.05) 28.23 (60.93) 25.51 (53.77) 26.22 (55.27) 26.98 (65.69)
Segment recall 32.05 (99.80) 29.71 (99.97) 20.14 (100.00) 35.63 (99.77) 35.74 (99.05) 32.95 (98.77) 33.48 (98.36) 33.38 (98.51) 31.64 (99.28)
Segment precision 34.05 (61.01) 4.00 (74.42) 1.55 (88.84) 23.95 (64.15) 1.14 (98.77) 9.36 (71.67) 7.04 (82.45) 15.85 (81.10) 12.12 (77.80)
Segment f1 31.39 (61.04) 6.91 (74.41) 2.77 (88.84) 26.95 (64.20) 2.19 (98.76) 14.31 (72.13) 11.30 (83.12) 20.98 (86.48) 14.60 (78.62)
Secondary Structure Element (SSE) Agreement Results. We ran a new expert agreement evaluation against the 8 basic SSEs (from DSSP). The summary is in 15. We do see above-random enrichment of SSEs in our tokens across all-metrics. The recall of existing SSEs is exceptional: $98.28\%$ block-level, $99.28\%$ segment-level, averaged over datasets and SSEs. This indicates the ability to recapitulate the 8 known elements, while the milder precision ($59.57$ block, $77.80$ segment) hints that GeoBPE goes beyond SSEs. As prior agreement results and case studies show, GeoBPE can find biologically meaningful regions (e.g. conserved homology, functional sites) and is not constrained to SSE boundaries, with the data dictating the exact high-level clusters.
@@ Appendix D Token Efficiency Metrics
Table 10: We evaluate token efficiency of GeoBPE across varying $|\mathcal{V}|\in\{600,2500,6000\}$, as reported in Figure 4, with Hyperparameter Settings 1.
@@
Table 11: We adopt the Small Structure Language Model evaluation protocol described in App. E. We sample 100 PDB structures. Best scTM and Designability scores are bolded. Underlined methods follow the evaluat -ion protocol in App. F.
@@
Method Codebook Size scTM Designability (scTM >0.5>0.5) Diversity (1 - mean TM) Uniqueness (TM=0.5)
@@ Appendix E Small Structure Language Model Evaluation
Goals & Aims. This section specifies a protocol used to compare the language modeling efficiency of GeoBPE vs VQ-VAE tokens. The goal of SSLM was to create a small, isolated environment for language model integration. Modeling is intentionally minimalistic–we train a small decoder-only Transformer (7.3M) parameters over the same PDB splits used in Sec. 4 (4848k structures). Fixing the same data splits, training, model architecture and hyperparameters (App. E.6), the aim of this experiment is to compare tokenizer options for a structure token language model. We emphasize the generation quality of the resulting models should be compared relatively; all models trained would be insufficient for real-world backbone design tasks. To bridge the gap with large-scale models, we provide an orthogonal study on scalability in App. F, where we reran SSLM with 1010x more data and model parameters.
Setup. For GeoBPE, we use the joint geometric vocabulary learned by GeoBPE; for VQ-VAE, we use their codebook. GeoBPE +SSLM incorporates the mask constraints used at generation time (Alg. 4) during training to ensure consistency between training and sampling. The same procedure is used for evaluating VQ-VAEs. For training and sampling, the only difference is dropping the mask constraints. For inference, the samples are passed through the VQ-VAE decoder to construct backbone coordinates instead of assembling the backbone directly via Alg. 5. This required considerably more resources, and we discuss how we implemented this in App. E.6.
@@
0: Corpus of proteins $\{t_{\tau}\}_{\tau=1}^{T}$ with final segmentations $\{\mathcal{P}^{(\tau)}\}$ and assigned medoids; joint vocabulary $\Sigma$ and tokenizers from Alg. 16, 17; a decoder-only Transformer $\mathsf{Tr}_{\theta}:\Sigma^{\ast}\!\to\!\Delta^{|\Sigma|}$ with causal mask; special BOS/EOS (optional); training steps $S$, optimizer $\mathcal{O}$.
@@
1:Dataset construction. For each $\tau$, build $x^{(\tau)}=\textsc{BackboneToSequence}(t_{\tau})$ (Alg. 17). Let $L_{\tau}=|x^{(\tau)}|$.
@@
0: Training corpus of tokenized backbones $\{x^{(\tau)}=(x^{(\tau)}_{1},\ldots,x^{(\tau)}_{L_{\tau}})\}_{\tau=1}^{T}$ constructed by Alg. 17 (motif, then $\theta,\omega,\phi$, repeating); valid sequence lengths satisfy $L_{\tau}\equiv 1\ (\mathrm{mod}\ 4)$ and end in a terminating motif token.
@@
0: Trained decoder-only Transformer $\mathsf{Tr}_{\theta}$ with vocabulary $\Sigma$ from Alg. 16; id blocks